EthanEnriquez:

2sin^2x-5sinx-3=0 Give a general formula of all the solutions

Ultrilliam:

@Vocaloid

Vocaloid:

start by factoring into: (2sin(x) + 1) * (sin(x) - 3) = 0 then solve 2sin(x) + 1 = 0 and sin(x) -3 = 0 use the unit circle if you are stuck

EthanEnriquez:

Giving a general for all solutions

EthanEnriquez:

Formula

Vocaloid:

Again, please solve: 2sin(x) + 1 = 0 and sin(x) -3 = 0 let me know if you are having trouble

EthanEnriquez:

You have to give one general formula

Vocaloid:

That's what I'm trying to do. Again, please solve: 2sin(x) + 1 = 0 and sin(x) -3 = 0 We do not give away direct answers here.